(1 point) Take the system x = 4x – xy, y = 5y + x2. How many critical points are there? What is the critical point with thePlease show work!!
Answer #1

. : 4-ay - y = 59 +32 G We know that Suppose Praound in an Rolated conical point of the systom. - fray) ay = gay) and f andOd (-2355,9) for (-ajj, 1) = 4 450 fy(-a55s, 4) = aus 3:03,) - Ox(-0355) - -4355 gyland) - 5 Jy (155, 1) = 5 Now matrix room